OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 14

OpenStudy (jiteshmeghwal9):

f(x)=-2x+8 & f(x)=14 so, -2x+8=14 solve for x :)

OpenStudy (anonymous):

how ? im pelletty at algebra

OpenStudy (anonymous):

sh1tty

OpenStudy (jiteshmeghwal9):

-2x=14-8 -2x=6 x=-3

OpenStudy (mayankdevnani):

f(14)=-2(14)+8=-28+8=-20

OpenStudy (mayankdevnani):

ok @kristopher101495

OpenStudy (jiteshmeghwal9):

hmmm.... Question is saying to find the value of x

OpenStudy (anonymous):

@mayankdevnani I think you are incorrect. it says f(x)=14 not f(14)

OpenStudy (jiteshmeghwal9):

@mayankdevnani

OpenStudy (anonymous):

how u get six from that ?

OpenStudy (jiteshmeghwal9):

u r incorrect :(

OpenStudy (mayankdevnani):

ok sry i got it

OpenStudy (mayankdevnani):

-2x+8=14 -2x=7/4 -x=7/4*2 x=-7/2

OpenStudy (jiteshmeghwal9):

since f(x)=-2x+8 & also 14 so -2x+8=14 now i subtract 8 from both sides & then divide -2 from both sides @kristopher101495 :)

OpenStudy (mayankdevnani):

f(-7/2)=-2(-7/2)+8=7+8=15

OpenStudy (mayankdevnani):

ok @kristopher101495

OpenStudy (anonymous):

ok i got it thanks

OpenStudy (mayankdevnani):

welcome